OpenStudy (anonymous):

Given f(x) = x^2 + 2, Find f(a+2) - f(a)

OpenStudy (alekos):

Have you attempted this?

OpenStudy (anonymous):

Yes.

OpenStudy (alekos):

what did you get?

OpenStudy (anonymous):

I wasn't sure how to get to the final answer, but f(a) = a^2 + 2; f(a+2)^2 +2. Correct?

OpenStudy (alekos):

f(a+2) = (a+2)^2 + 2

OpenStudy (anonymous):

Yes. I missed the first part f(a+2). But how should I proceed?

OpenStudy (alekos):

well f(a+2) - f(a) = (a+2)^2 +2 - (a^2+2) and go from there

OpenStudy (anonymous):

@alekos my final answer was 8 correct?

OpenStudy (alekos):

No. Following on we get.... (a^2+4a+4)+2 - a^2 - 2 try the next step

OpenStudy (anonymous):

4 + 4a..yeah ..he is guiding u correct ..:)

OpenStudy (anonymous):

My final answer was 4a + 4

OpenStudy (alekos):

Yep, that's it! well done

OpenStudy (anonymous):

@alekos Thank you. You're a lifesaver!

OpenStudy (alekos):

No problem, you're most welcome